### v1

# SFMNet: Sparse Focal Modulation for 3D Object Detection

Abstract
We propose SFMNet, a novel 3D sparse detector that combines the efficiency of sparse convolutions with the ability to model long-range dependencies. While traditional sparse convolution techniques efficiently capture local structures, they struggle with modeling long-range relationships. However, capturing long-range dependencies is fundamental for 3D object detection. In contrast, transformers are designed to capture these long-range dependencies through attention mechanisms. But, they come with high computational costs, due to their quadratic query-key-value interactions. Furthermore, directly applying attention to non-empty voxels is inefficient due to the sparse nature of 3D scenes. Our SFMNet is built on a novel Sparse Focal Modulation (SFM) module, which integrates short- and long-range contexts with linear complexity by leveraging a new hierarchical sparse convolution design. This approach enables SFMNet to achieve high detection performance with improved efficiency, making it well-suited for large-scale LiDAR scenes. We show that our detector achieves state-of-the-art performance on autonomous driving datasets.
1 Introduction
3D object detection has become fundamental due to its potential applications in autonomous driving [17, 2, 41, 29, 3, 50], and robotics [44, 66]. Extensive research has explored various methods, including point-based [32, 33, 34, 35, 36, 38, 39, 54], range-based [13, 42, 45], and voxel-based approaches [63, 5, 10, 23, 37, 52, 55, 56, 61, 62, 7, 47, 59, 58, 40]. Point-based methods process raw point clouds but often suffer from high computational costs or reduced accuracy in large-scale scenes due to point sampling. Range-based methods project 3D points to apply 2D convolutions, making them computationally efficient. However, they often struggle with occlusions, object localization, and size regression errors, as they lack the expressive 3D features necessary for accurate detection.
Refer to caption(a) Sparse convolution detectors (e.g., CenterPoint)3D backbone2D backboneBEVDetection head(b) Transformer detectors (e.g., DSVT)3D backbone2D backboneBEVDetection headRegional GroupingRegional GroupingRegional GroupingRegional Grouping(c) SFMNet (ours)3D backbone2D backboneBEVSparse detection headDenseconv.blockSparseconv.blockTransf.moduleSFM.module
Figure 1: SFMNet in comparison with other architectures. Similar to sparse convolution-based detectors, e.g.,[56] (a), our SFMNet detector (c) comprises a multi-stage 3D backbone with sparse convolution blocks. In contrast, SFMNet is fully sparse, including the 2D backbone, and integrates a novel SFM module (light orange) into both the 3D and 2D backbones. Furthermore, unlike transformers, SFMNet uses sparse convolutions as the token mixer, which makes query modulation efficient. Moreover, unlike the single-stride network design used in transformer-based detectors (e.g., [47]), our SFMNet detector employs an encoder-based design.
Voxel-based methods, which are the focus of this paper, convert the irregular point cloud into regular voxels. This allows the use of convolutions, specifically sparse convolutions, which operate efficiently in the presence of empty voxels [52, 20, 9]. However, while efficient at capturing local structures, methods built on submanifold sparse convolution [19] significantly reduce the receptive field, limiting the representation capacity and long-range dependencies that are crucial for deep neural networks [49]. In 3D detection, long-range dependencies refer to connections between voxels within the same object and its immediate surroundings rather than between distant objects. Since LiDAR data is inherently sparse and occlusions are common, relying solely on local features makes it nearly impossible to detect objects like pedestrians with only a few sampled points. However, contextual cues, such as a crosswalk, can significantly aid detection, as even a few points above it may suggest the presence of a pedestrian. Therefore, long-range dependencies are essential for capturing contextual information.
To address these limitations, the use of transformers has been proposed [46, 14, 47, 65, 43, 8, 48]. With self-attention (SA) as their core mechanism, transformers can capture global relationships, increasing the detector representation capacity. Despite its benefits, the quadratic complexity of SA is a significant bottleneck, particularly in autonomous vehicle scenarios where LiDAR scenes span hundreds of meters and contain hundreds of thousands of points.
On the other hand, convolutional neural networks have recently made a comeback, primarily in 2D, through the use of large kernels [12, 11, 25]. Large-kernel convolution-based methods improve representation capacity by increasing the receptive field and offering greater robustness compared to transformers [4]. However, this approach is often impractical for 3D scenes, as the number of parameters grows cubically, necessitating techniques like sophisticated kernel weight partitioning with shared weights [6], data-driven kernel generation [27], or weight pruning [16]. Although these methods reduce parameter bottlenecks and enlarge the effective receptive field, they offer minimal accuracy improvements and are less efficient than traditional sparse convolutions.
Inspired by these solutions, this paper aims to combine the best of both worlds—enabling transformer-like long-range dependencies while maintaining the efficiency of sparse convolutions. We introduce a new 3D sparse detector, termed SFMNet, that relies entirely on (sparse) convolutions; however, between these convolutions, we introduce a novel ’transformer-like’ module where the attention mechanism is also replaced with convolutions. Our SFM module captures both short- and long-range relationships between voxels, bridging gaps left by traditional sparse convolution detectors while providing greater efficiency than transformer-based detectors. As illustrated in Fig. 1, our model differs from traditional sparse convolution detectors by being fully sparse, including its 2D backbone, and by incorporating our SFM module within the backbones.
The Sparse Focal Modulation (SFM) module modulates short- and long-range contexts and integrates them with query tokens, similar to transformers. However, instead of the quadratic complexity of self-attention, SFM operates with linear complexity relative to the number of tokens (voxels). The core idea is to efficiently extract local contexts using sparse convolutions in a hierarchical structure, enabling the design to capture global relationships. Limiting attention to local context windows has been shown to reduce the complexity bottleneck [26]. However, due to the sparse nature of LiDAR points, the number of occupied voxels in each local window can vary significantly, making efficient local-based attention non-trivial [47]. To efficiently extract context windows, we design our SFM module with stacked submanifold sparse convolutions, using small kernels and varying dilation rates. This design enables hierarchical context aggregation, progressively expanding the receptive field at each level. Sparse convolutions capture local structures efficiently, while the hierarchical design and increased dilation rates allow for learning sparse long-range dependencies.
We validate the effectiveness of our detector on three well-known autonomous driving datasets. Our results demonstrate that the network advances the state of the art in long-range detection (Argoverse2 [50]), for which it is designed. On mid- to short-range datasets, namely Waymo Open [41] and nuScenes [2], it performs comparably or slightly better than previous state-of-the-art detectors.
Thus, the main contributions are twofold:
  • We propose a novel sparse focal modulation (SFM) module. It enables long-term dependencies, similar to transformers, while maintaining efficiency like sparse convolutions. This is achieved by hierarchically aggregating contextual information from sparse inputs. The SFM module outperforms traditional attention-based methods by reducing computational overhead and efficiently processing sparse voxels of varying densities within context windows.
  • We introduce SFMNet, a 3D object detector built on top of the SFM module, which effectively handles large point cloud scenes. We validate the detector’s usefulness on well-known large-scale LiDAR datasets, achieving state-of-the-art results.
2 Related work
3D detection on point clouds. 3D detection methods can be broadly categorized into point-based methods and voxel-based methods. Point-based methods use PointNet [32, 33]-like architectures to extract geometric features directly from raw point clouds [32, 33, 34, 35, 36, 38, 39, 54]. However, in large-scale point cloud scenes typical of outdoor LiDAR datasets in autonomous driving, these methods suffer from lower accuracy and high processing times due to point sampling and neighbor searching. In contrast, voxel-based methods convert irregular point clouds into regular voxel grids and apply sparse convolutions on non-empty voxels [63, 5, 10, 23, 37, 52, 55, 56, 61, 62, 7, 59, 58, 40]. While efficient for processing large-scale data, these methods rely on submanifold sparse convolution, which might limit network representation capacity by prioritizing efficiency over capturing long-range dependencies.
Long-range dependencies for 3D detection. Recent 3D detection methods capture long-range dependencies using large-kernel sparse convolutions or transformers. Large-kernel convolution approaches expand the receptive field by increasing kernel sizes, a technique that has shown accuracy gains in 2D detection [12, 11, 25]. However, directly applying large kernels to 3D scenes incurs a cubically increasing computational overhead. To address this, LargeKernel3D [6] introduces a spatial partitioning convolution with shared weights among neighboring voxels. LinK [27] further optimizes efficiency by generating sparse kernels dynamically, assigning weights only to non-empty voxels rather than maintaining a dense kernel matrix. Similarly, LSK3DNet [16] reduces computation by pruning redundant weights across spatial and channel dimensions. While these methods alleviate some computational demands, they still face performance challenges in large-scale 3D data.
Transformer-based detectors, by contrast, use self-attention to capture long-range dependencies, enabling global token interactions. However, applying attention in high-resolution 3D scenes with many tokens is computationally prohibitive [65, 8, 48]. To mitigate this, VoTr [30] implements local and dilated attention, focusing on sparse voxel interactions. SWFormer [43] utilizes a window-based transformer with a bucketing scheme to manage windows with varying numbers of non-empty voxels. SST [14] refines token processing by grouping tokens regionally, applying attention within each group, and shifting regions to capture structured interactions. DSVT [47] sequentially applies windowed self-attention along the X and Y axes to capture dependencies with reduced complexity. Although they improve long-range relationships, they still face quadratic scaling issues with context window tokens due to self-attention.
We propose a network that combines the strengths of sparse convolutions and transformers. Our design maintains the efficiency of sparse convolutions while enabling long-range dependencies similar to transformers. However, unlike traditional transformers, our module is built from sparse convolutions and operates with linear complexity relative to the number of tokens (voxels). Furthermore, eliminating the need for sampling helps us achieve better accuracy.
[…]
Figure 2: SFM module vs. Local attention. Given a query token (burnt orange), local-based attention (a) identifies and groups non-empty tokens (voxels) within a context window (light gray) to form key and value tokens. The query-key-value interactions have quadratic complexity. In contrast, SFM (b) leverages sparse convolutions to extract multiple levels of context, reducing query interactions to only a few focal contexts and achieving linear complexity with respect to the number of non-empty voxels in the context window. Furthermore, while local-based attention suffers from inefficiencies in batch processing due to the varying numbers of non-empty voxels within each context window, our SFM approach is robust to this variability, extracting a constant number of contexts for each window.
3 Method
Voxel-based detectors are widely used for 3D detection in outdoor scenes, such as those in autonomous driving scenarios. They commonly leverage sparse convolutions for efficient learning in sparse environments. There are two main types of sparse convolutions: regular sparse convolution [18] and submanifold sparse convolution [19]. Regular sparse convolutions expand features into (not necessarily occupied) neighboring voxels, enabling information exchange between spatially disconnected regions, but increasing feature map density. Due to their expansive nature, they are typically used for downsampling and upsampling. Submanifold sparse convolutions, on the other hand, operate strictly on occupied voxel features, ensuring that the output retains the same sparsity as the input, thereby facilitating efficient learning of local structures. However, due to their limited kernel size, they lack long-range dependencies, which are useful for object detection [49].
We aim to design an efficient detector built from sparse convolutions capable of capturing both short- and long-range dependencies. A straightforward approach to modeling global relationships would be to use transformers, which have shown great success in the 2D domain. However, applying standard transformers to sparse point clouds is challenging: (1) With numerous occupied voxels in each scene, standard self-attention suffers from massive query-key and query-value interactions, leading to a high computational overhead or suboptimal accuracy due to subsampling [31, 60]. (2) Due to the sparse nature of LiDAR point clouds, the number of occupied voxels in each local region can vary significantly, leading to inefficient batch processing [14, 47].
To address both the computational overhead and the varying number of occupied voxels within a context window, we propose replacing self-attention with Sparse Focal Modulation (SFM) as the token mixer in the transformer. The SFM module reduces query interactions to only a few focal contexts around each query by hierarchically extracting representations of the surrounding connections. Additionally, sparse convolutions were designed to efficiently handle varying numbers of occupied voxels within each local region and can facilitate effective context extraction.
The context extraction and interaction phases of our proposed SFM module are illustrated in Fig. 2, in comparison to naive local-based attention. Local-based attention involves a context extraction phase for each query token, followed by query-key and query-value interactions. The context extraction phase requires identifying all non-empty voxels within the context window, sampling or padding them to a fixed maximum number of neighbors, and concatenating their features for batch processing. In high-resolution scenes with hundreds of thousands of voxels, this process quickly becomes a bottleneck. Moreover, the query-key-value interactions result in quadratic complexity relative to the number of tokens within the context window. In contrast, SFM modulates a query’s interactions by extracting multiple levels of context using sparse convolutions, focusing on short-range structures at lower levels and long-range dependencies at higher levels. The sparse convolutions are designed to handle varying numbers of non-empty voxels efficiently within different context windows, enabling effective batch processing. By using (sparse) convolutions with different receptive fields for context extraction, the interaction phase is reduced to only the number of extracted context levels. Our SFM module is described in Section 3.1.
The effectiveness of our proposed SFM is demonstrated through a novel sparse detector, called SFMNet, introduced in Section 3.2. Our detector effectively captures long-range dependencies with a large receptive field, improving performance in object detection tasks.
3.1 Sparse focal modulation
Focal Modulation (FM) is an alternative token mixer to self-attention, designed to limit token interactions. It has been applied in 2D tasks, such as classification, segmentation, and detection [53, 28]. Extending FM to 3D faces two primary challenges. First, 3D representations are inherently sparse. The voxel count, including both empty and non-empty voxels, scales cubically. For example, in Argoverse2 [50], a typical scene covers an area of $400\times 400\times 8$ meters, resulting in approximately $640$ million voxels at a resolution of $0.1\times 0.1\times 0.2$ meters. With only about 100,000 sampled points, this results in a high percentage of empty voxels. The second challenge is that 2D FM typically relies on large kernels to capture global context by using a stack of depth-wise convolutional layers with progressively larger kernels to capture visual contexts at various scales. However, in 3D object detection, sparse depth-wise convolutions have shown no advantage over spatial convolutions and are memory-intensive [6]. Furthermore, applying large kernels in 3D scenes leads to a cubic increase in parameter count, substantially raising computational costs.
We propose a new Sparse Focal Modulation (SFM) module that addresses these challenges. It captures both short- and long-range contexts while adaptively exchanging information with each token query. Unlike the conventional approach of using large kernels to expand the context window densely, we expand the context window sparsely with small kernels and increased dilation rates. We reason that queries related to small objects often rely on local structures and benefit from compact, focused context windows. In contrast, queries for large objects (e.g., partially occluded vehicles or roads) require broader context windows, but not necessarily dense windows, given the inherent sparsity of LiDAR point clouds. Hence, our context aggregation is sparse.
Fig. 3 illustrates our SFM module, which modulates each non-empty voxel feature ($x$) with surrounding information to produce a spatially aware feature ($z$). Its core component, the Context Aggregation (CA) block, captures relevant context for each query token by: 1) efficiently extracting contexts hierarchically, from local to global regions (e.g., red, green, and blue regions in Fig. 2); and 2) adaptively aggregating context features at each level using a gating mechanism to provide meaningful context for each query. The resulting aggregated context of the CA block, ctx, is modulated with the corresponding query $q$ to generate the feature $z$.
[…]
Figure 3: Sparse Focal Modulation (SFM). Given non-empty voxel features $x$, the query $q$ and initial focal features $f^{0}$ are generated through linear layers. $f^{0}$ is fed into the Context Aggregation module, which aggregates multi-level contexts by stacking submanifold sparse convolutions (SubMConv) with small kernel sizes ($k$) and varying dilation rates ($d$), each followed by a GeLU activation. Context aggregation is achieved through a gated mechanism that adaptively combines features from each level ($f^{\ell}$) based on their relevance. The final context (ctx) is mapped to the query space using a linear layer $h(\cdot)$ and then modulated with the query $q$ to produce spatially-aware features $z$. This design captures both local and global dependencies while maintaining computational efficiency.
Submanifold sparse convolutions efficiently handle voxel sparsity, as they operate strictly on non-empty voxels. By stacking submanifold sparse convolutions with small kernels, we can effectively capture local contexts at each level. Specifically, given a feature vector $x\in\mathbb{R}^{N\times C}$, where $N$ represents the number of voxels and $C$ denotes the number of features, we project this vector into new feature spaces using linear layers to generate both per-voxel queries $q\in\mathbb{R}^{N\times C}$ and initial focal features $f^{0}\in\mathbb{R}^{N\times C}$. The focal features are then fed into the Context Aggregation (CA) module. The CA produces $L$ levels of focal feature maps, $f^{\ell}$, with each level fed to a submanifold sparse filter that extracts the next focal context. Each level has a progressively larger effective receptive field than the previous level. At level $\ell$, the focal feature $f^{\ell}\in\mathbb{R}^{N\times C}$ is computed as
$f^{\ell}=\text{GeLU}(\text{SubMConv}(f^{\ell-1};k^{\ell},d^{\ell})),$ (1)
where $\text{SubMConv}(\cdot;k^{\ell},d^{\ell})$ denotes a submanifold sparse convolution with kernel size $k^{\ell}$ and dilation rate $d^{\ell}$, and GeLU is an activation function [21]. The hierarchical convolutions effectively capture multiple levels of focal abstraction, with the effective receptive field at level $\ell$ given by:
$r^{\ell}=1+\sum_{i=1}^{\ell}(k^{i}-1)\cdot d^{i}.$ (2)
As for the second key principle, when aggregating context features, we note that not all queries benefit equally from short-range or long-range contexts. Specifically, queries for small objects are more likely to gain from the short-range context available at lower focal levels, while larger objects or background regions tend to benefit more from the long-range context captured at higher focal levels. To address this, a gating mechanism is employed to weigh the focal features from the different levels. This gate adaptively regulates the amount of context extracted from each level, enabling the CA module to learn suitable contexts for various queries. By combining the focal features with the gate weights, the CA produces a context feature vector for each query, defined as:
$\text{ctx}=\text{h}(\sum_{\ell=1}^{L}f^{\ell}\cdot g^{\ell})\in\mathbb{R}^{N\times C},$ (3)
where $g^{\ell}\in\mathbb{R}^{N\times 1}$ represents the gate weight for focal level $\ell$, and $h(\cdot)$ is a linear layer that projects the output context into the query space.
Finally, we modulate the focal contexts to the queries through element-wise multiplication:
$z=q\cdot\text{ctx}\in\mathbb{R}^{N\times C}.$ (4)
This enables each query to learn its relevant context.
In summary, our module addresses the aforementioned challenges. By stacking sparse convolutions with small kernels and progressively increasing dilation, we effectively aggregate multi-level contexts. Efficiently handling voxel sparsity while modulating a query with its aggregated context maintains token interactions at linear complexity.
Method mAPH Vehicle Pedestrian Cyclist
Baseline 66.9 64.4 67.0 69.1
SFMNet 69.5 68.1 69.2 71.1
Table 1: The utility of our SFM module. Incorporating a single SFM module into each backbone (3D and 2D) within CenterPoint improves L2 mAPH by $2.6\%$ on a $20\%$ Waymo Open subset.
To demonstrate the utility of our proposed module, we incorporated a single SFM module into each backbone (3D and 2D) within the CenterPoint [56] architecture, without otherwise altering the structure. As shown in Table 1, this addition yielded a $2.6\%$ improvement in the Level-2 mAPH on a subset of the Waymo Open dataset [41]. Next, we will introduce our architecture that leverages our module.
3.2 SFMNet
Building on our sparse focal modulation (SFM) module, we present SFMNet, a simple and efficient sparse detector for 3D object detection. It enables learning long-range dependencies that convolution-based backbones lack, while maintaining sparse convolution efficiency.
Our architecture is built on two core principles. First, increasing the effective receptive field (ERF) in mid- and high-level layers enhances the network’s capacity to capture broader spatial contexts. This is achieved by incorporating our proposed SFM module. Second, once a reasonable receptive field is achieved, scaling a model with small kernels becomes more efficient due to better parameter utilization. This aligns with [12]’s observation that beyond a certain point, adding large kernels is less efficient than adding small kernels. In our setup, SFM blocks can be seen as analogous to large kernels, while SRB blocks correspond to small kernels. Therefore, we include more SRB blocks than SFM blocks (see Section 4.2).
[…]
Figure 4: SFMNet architecture. (a) The SFMNet architecture extracts voxel features from a point cloud input (VFE), processes them through multiple stages, and converts them to a bird’s-eye view (BEV) representation before passing the output to a 2D backbone and detection head. (b) Each stage consists of multiple Sparse Focal Modulation (SFM) blocks followed by Sparse Residual Blocks (SRBs). (c) The SFM block aggregates contextual information and processes it, while the SRB employs SubMConv layers with batch norm (BN) and ReLU activations to capture spatial features.
Fig. 4(a) illustrates the SFMNet architecture. The input point cloud is voxelized, and features are extracted using a dynamic voxel feature extraction (VFE) method [64]. These features are then processed by our 3D and 2D sparse backbones. Our 3D backbone includes four stages and between each pair of consecutive stages, a regular sparse convolution with stride $2$ is used for downsampling. Each stage includes an SFM block followed by multiple submanifold residual blocks (SRBs) for high-level feature extraction (Fig. 4(b)). We use a variable number of SFM blocks and SRBs to enhance representational capacity and increase depth depending on the dataset. The higher the stage, the more SRB blocks per SFM block should be added to scale the model.
The SFM block utilizes our SFM module as the token mixer within a standard MetaFormer-like design [57], as shown in Fig. 4(c). Formally, given an input feature map $x\in\mathbb{R}^{N\times C}$ and its corresponding focal-modulated feature $z\in\mathbb{R}^{N\times C}$ generated by the SFM module, as described in Section 3.1, the SFM block is formulated as:
$y^{\prime}$ $=\text{LN}(z)+x$ (5)
$y$ $=\text{LN}(\text{MLP}(y^{\prime}))+y^{\prime},$ (6)
where LN denotes LayerNorm, and the multilayer perceptron (MLP) contains a single hidden layer. The 3D sparse features are then compressed into 2D sparse bird’s-eye view (BEV) features, following the approach in [58], and subsequently fed into a 2D backbone. Similar to the 3D backbone, the 2D backbone also consists of an SFM block followed by SRBs but employs 2D sparse convolutions instead of 3D.
4 Experiments
Datasets and metrics. To validate our approach’s effectiveness, we conducted experiments on three popular datasets: Argoverse2 [50], Waymo Open [41], and nuScenes [2]. The detection range of these datasets varies, with long-range detection in Argoverse2 (up to ±$\pm 200$ meters), mid-range in Waymo Open (up to ±$\pm 75$ meters), and short-range in nuScenes (up to ±$\pm 54$ meters). We use the standard evaluation metrics for each dataset: mAP for Argoverse2, mAP and mAP weighted by heading accuracy (mAPH) for Waymo Open, and mAP and nuScenes detection score (NDS) for nuScenes. On Waymo, the metrics are applied to two difficulty levels: L1 for objects with more than five LiDAR points and L2 for those with at least one sampled point.
Method
mAP
Vehicle
Bus
Pedestrian
Stop Sign
Box Truck
Bollard
C-Barrel
Motorcyclist
MPC-Sign
Motorcycle
Bicycle
A-Bus
School Bus
Truck Cab
C-Cone
V-Trailer
Sign
Large Vehicle
Stroller
Bicyclist
Truck
MBT
Dog
Wheelchair
W-Device
W-Ride
CenterPoint [56] 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
VoxelNeXt [7] 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
HEDNet [59] 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
SAFDNet [58] 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
SFMNet (Ours) 40.4 79.0 49.5 72.3 49.3 46.1 65.1 73.8 55.1 53.2 62.8 54.2 30.2 44.8 26.7 57.9 31.7 20.9 8.3 33.4 44.4 21.6 0.0 22.1 3.1 33.1 12.1
Table 2: Results on the Argoverse2 validation set (long-range, up to ±$\pm 200$ meters). The long-range detection dataset highlights our detector’s ability to capture long-range dependencies.
Design Method mAP/mAPH Vehicle AP/APH Pedestrian AP/APH Cyclist AP/APH
L1 L2 L1 L2 L1 L2 L1 L2
Transformers SST [14] [CVPR’22] 74.5/71.0 67.8/64.6 74.2/73.8 65.5/65.1 78.7/69.6 70.0/61.7 70.7/69.6 68.0/66.9
TransFusion-L [1] [CVPR’22] -/- -/64.9 -/- -/65.1 -/- -/63.7 -/- -/65.9
SWFormer [43] [ECCV’22] -/- -/- 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 -/- -/-
CenterFormer [65] [ECCV’22] 75.6/73.2 71.4/69.1 75.0/74.4 69.9/69.4 78.0/72.4 73.1/67.7 73.8/72.7 71.3/70.2
DSVT-Voxel [47] [CVPR’23] 80.3/78.2 74.0/72.1 79.7/79.3 71.4/71.0 83.7/78.9 76.1/71.5 77.5/76.5 74.6/73.7
PTv3 [51] [CVPR’24] -/- 73.0/70.5 -/- 71.2/70.8 -/- 76.3/70.4 -/- 71.5/70.4
Large kernels LargeKernel3D [6] [CVPR’23] -/- -/- 78.1/77.6 69.8/69.4 -/- -/- -/- -/-
LinK [27] [CVPR’23] (6 epochs) -/- -/64.6 -/- -/65.0 -/- -/60.4 -/- -/68.4
Small kernels SECOND [52] [Sensors’18] 67.2/63.1 61.0/57.2 72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0
PointPillars [23] [CVPR’19] 69.0/63.5 62.8/57.8 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9
Part-A2 [37] [TPAMI’20] 73.6/70.3 66.9/63.8 77.1/76.5 68.5/68.0 75.2/66.9 66.2/58.6 68.6/67.4 66.1/64.9
PV-RCNN [36] [CVPR’20] 76.2/73.6 69.6/67.2 78.0/77.5 69.4/69.0 79.2/73.0 70.4/64.7 71.5/70.3 69.0/67.8
CenterPoint [56] [CVPR’21] 75.9/73.5 69.8/67.6 76.6/76.0 68.9/68.4 79.0/73.4 71.0/65.8 72.1/71.0 69.5/68.5
AFDetV2 [22] [AAAI’22] 77.2/74.8 71.0/68.8 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1
FSD [15] [NeurIPS’22] 79.6/77.4 72.9/70.8 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3
PV-RCNN++ [38] [IJCV’23] 78.1/75.9 71.7/69.5 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2
VoxelNeXt [7] [CVPR’23] 78.6/76.3 72.2/70.1 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2
HEDNet [59] [NeurIPS’23] 81.4/79.4 75.3/73.4 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9
SAFDNet [58] [CVPR’24] 81.8/79.9 75.7/73.9 80.6/80.1 72.7/72.3 84.7/80.4 77.3/73.1 80.0/79.0 77.2.76.2
SFMNet (ours) 81.8/79.8 75.7/73.8 80.5/80.0 72.6/72.2 85.0/80.6 77.4/73.2 79.9/78.9 77.0/76.1
Table 3: Results on the Waymo Open validation set (mid-range, up to ±$\pm 75$ meters). Our SFMNet outperforms transformer- and large-kernel-based detectors, while achieving results comparable to the state-of-the-art small-kernel detector [58].
Implementations details. To compare with prior state-of-the-art methods, we trained SFMNet for $24$ epochs on Argoverse2 and Waymo Open, and $20$ epochs on nuScenes. Ablation studies used 6 epochs on Argoverse2. Further implementation details are in the supplemental materials.
4.1 Results
Quantitative results. We expect our model to be particularly beneficial for long-range scenes. Indeed, as shown in Table 2, which compares our method to state-of-the-art sparse convolution-based detectors with small kernels, our detector achieves a $0.7\%$ mAP improvement over the previous best method [58], demonstrating its ability to capture long-range dependencies. For mid-range detection on Waymo Open, Table 3 compares our method to state-of-the-art methods on the validation set, where ’-’ indicates unavailable results. Our detector outperforms both transformer-based and large kernel-based methods, achieving gains of $1.7\%$ and $3.3\%$ mAPH on Level-2 difficulty compared to the transformer-based detectors DSVT [47] and PTv3 [51], respectively, while also achieving comparable results to the best small kernel-based detector [58]. For the short-range detection dataset nuScenes,Table 4 shows that our method achieves SoTA results on the mAP metric and ranks second on the NDS metric, outperforming both large kernel and transformer-based methods. The effective receptive field (ERF) is $3.3$ meters for Argoverse2 and $2$ meters for Waymo Open and nuScenes.
Refer to caption Refer to caption Refer to caption
(a) Input scene (b) ERF without SFM (c) ERF with SFM
Figure 5: Effective receptive field (ERF) of SFMNet. (a) A typical input scene with height color-coded in shades of gray. (b,c) The ERF for each ground-truth object, color-coded from blue (outside ERF) to red (high attention regions). Red and orange rectangles highlight zoomed-in views of two regions. In the right (orange) zoom-in, four stationary cars and a van, plus two moving cars and a van, are shown. Without the SFM module (b), only a local portion of each vehicle is within the ERF. With the SFM module (c), the entire object is captured, providing more context and spatial understanding for accurate 3D detection.
Qualitative results. To qualitatively validate our detector’s ability to capture an extensive ERF, for each ground-truth object, we calculate the ERF by sampling a query voxel and analyzing its reach. Fig. 5 visualizes the ERF (color-coded from blue to red) of our detector with and without our SFM module. The more yellowish the color, the larger the ERF. For each ground-truth object, we randomly select one voxel to serve as the query and calculate the ERF associated with it. As intended, the dependency coverage is regional, resulting in large regional receptive fields. Without the SFM module (b), the detector’s receptive field is limited to local areas around each query, capturing only fragments of the objects. For instance, in the left zoom-in (red rectangle) in Fig. 5(b), there are four vehicles. The receptive field covers only portions of each vehicle, leaving most voxels outside this field (blue). In contrast, with our SFM module (c), the detector gains a broader receptive field, allowing it to capture even partially occluded vehicles. This is reflected in the four vehicles being highlighted in green, yellow, and red colors.
Design Method NDS mAP Car Truck Bus Trailer C.V. Pedestrian Motor Bike Cone Barrier
Transformers TransFusion-L [1] [CVPR’22] 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
DSVT-Voxel [47] [CVPR’23] 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.8 70.9
Large kernels LargeKernel3D [6] [CVPR’23] 69.1 63.9 85.1 60.1 72.6 41.4 24.3 85.6 70.8 59.2 72.3 67.7
LinK [27] [CVPR’23] 69.5 63.6 - - - - - - - - - -
Small kernels CenterPoint [56] [CVPR’21] 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
VoxelNeXt [7] [CVPR’23] 66.7 60.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
PillarNeXt-L [24] [CVPR’23] 68.4 62.2 85.0 57.4 67.6 35.6 20.6 86.8 68.6 53.1 77.3 69.7
HEDNet [59] [NeurIPS’23] 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
SAFDNet [58] [CVPR’24] 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
SFMNet (ours) 71.3 66.9 87.6 63.2 76.2 48.9 26.0 88.4 74.7 56.9 78.5 69.1
Table 4: Results on the nuScenes validation set (short-range, up to ±$\pm 54$ meters). Our SFMNet outperforms all transformer- and large-kernel-based detectors, surpassing the previous SoTA small-kernel detector on the mAP metric and ranking second on the NDS metric.
4.2 Ablation study
Table 5 compares our SFM module to traditional global and local attention, as well as set attention [47], on the Argoverse2 dataset over six epochs. For this evaluation, we replaced an attention block in the 3D and 2D backbones with either a single SFM module or an alternative attention mechanism, demonstrating that our SFM achieves higher accuracy and efficiency. While global attention attains a high ERF, it quickly reaches memory limits ($80$ GB) due to the large voxel count. In contrast, local and set attention operate within a limited context window, with set attention outperforming local attention. However, our SFM module surpasses both in accuracy (mAP) and efficiency (memory usage) despite using a smaller context window.
Method Window size(m) Memory(GB) mAP
Global-attention 400 OOM -
Local-attention 3.3 44.06 34.8
Set-attention (DSVT [47]) 3.6 44.21 35.5
SFM (ours) 1.9 32.76 36.8
SFM (ours) 3.3 37.53 37.1
Table 5: Comparison to attention mechanism. We replace our SFM module with various attention mechanisms and report mAP and per-GPU memory peak usage. While global attention covers the entire ERF, it quickly exceeds memory limits. In contrast, Local-attention and Set-attention [47] restrict their focus to voxels within a local context window. Our SFM module outperforms both in accuracy and efficiency, even with a smaller context window.
Table 6 examines the effect of varying the number of SFM blocks and SRBs. It shows that increasing the number of SRBs after the SFM block in Stage 4 (A) improves accuracy with a modest parameter increase. Similarly, adding more SFM block repetitions with SRBs (B) boosts accuracy but at a higher parameter cost. A balanced approach, with moderate repetitions and additional SRBs (C), achieves a good trade-off between accuracy and parameters.
More ablation studies on SFM module placement and ERF size choices are provided in the supplemental materials.
Stage 4 2D backbone #Param.(M) mAP
#SRB Multiplier #SRB Multiplier
A 2 1 4 1 3.84 36.7
4 1 4 1 4.29 37.2
6 1 4 1 4.73 37.5
B 2 2 4 2 5.27 38.0
2 4 4 2 7.14 38.6
2 6 4 2 9.01 39.1
C 4 2 4 2 6.16 38.9
Table 6: Effect of SFM blocks and SRBs. Increasing SRBs after each SFM block in Stage 4 and the 2D backbone boosts accuracy. Adding more SFM block repetitions (Multiplier) further improves accuracy, though at the cost of more parameters. A balanced trade-off between accuracy and parameters is achieved in (C).
Limitations. SFMNet relies on sparse convolutions with custom CUDA operations, making it less deployment-friendly than other deep learning frameworks (e.g., PyTorch).
5 Conclusion
This paper introduced a novel 3D Focal Modulation module, Sparse Focal Modulation (SFM), designed to effectively capture dependencies from short- to long-range. It is based on a hierarchical submanifold sparse convolution design to efficiently handle data sparsity. It also presented SFMNet, a new sparse convolution-based 3D detector built on top of the SFM. It addresses the limitation of traditional (sparse) convolution detectors by significantly enlarging the effective receptive field. Experimental results demonstrate that SFMNet achieves state-of-the-art performance on long-range detection while delivering comparable or superior results on mid- to short-range datasets.

### v2

@@
 While traditional sparse convolution techniques efficiently capture local structures, they struggle with modeling long-range relationships.
-However, capturing long-range dependencies is fundamental for 3D object detection.
+However, these relationships are essential for 3D object detection.
 In contrast, transformers are designed to capture these long-range dependencies through attention mechanisms.
@@
 1 Introduction
-3D object detection has become fundamental due to its potential applications in autonomous driving [17, 2, 41, 29, 3, 50], and robotics [44, 66].
-Extensive research has explored various methods, including point-based [32, 33, 34, 35, 36, 38, 39, 54], range-based [13, 42, 45], and voxel-based approaches [63, 5, 10, 23, 37, 52, 55, 56, 61, 62, 7, 47, 59, 58, 40].
+3D object detection is fundamental for applications in autonomous driving [nuscenes, chang2019argoverse, kitti1, waymo, once, wilson2023argoverse] and robotics [thrun2006stanley, zhu2017target].
+Extensive research has explored various methods, including point-based [qi2017pointnet, qi2017pointnet++, qi2018frustum, pointrcnn, pvrcnn, pvrcnn++, shi2020point, yang20203dssd], range-based [fan2021rangedet, sun2021rsn, tian2022fully], and voxel-based approaches [chen2019fast, chen2023voxelnext, voxelrcnn, pointpillars, parta2, shrout2023gravos, wang2023dsvt, second, ye2020hvnet, yin2021center, zhang2024safdnet, zhang2024hednet, zheng2021cia, zheng2021se, zhou2018voxelnet, song2023vp].
 Point-based methods process raw point clouds but often suffer from high computational costs or reduced accuracy in large-scale scenes due to point sampling.
@@
 Figure 1: SFMNet in comparison with other architectures.
-Similar to sparse convolution-based detectors, e.g.,[56] (a), our SFMNet detector (c) comprises a multi-stage 3D backbone with sparse convolution blocks.
+Similar to sparse convolution-based detectors, e.g.,[yin2021center] (a), our SFMNet detector (c) comprises a multi-stage 3D backbone with sparse convolution blocks.
 In contrast, SFMNet is fully sparse, including the 2D backbone, and integrates a novel SFM module (light orange) into both the 3D and 2D backbones.
 Furthermore, unlike transformers, SFMNet uses sparse convolutions as the token mixer, which makes query modulation efficient.
-Moreover, unlike the single-stride network design used in transformer-based detectors (e.g., [47]), our SFMNet detector employs an encoder-based design.
-Voxel-based methods, which are the focus of this paper, convert the irregular point cloud into regular voxels.
-This allows the use of convolutions, specifically sparse convolutions, which operate efficiently in the presence of empty voxels [52, 20, 9].
-However, while efficient at capturing local structures, methods built on submanifold sparse convolution [19] significantly reduce the receptive field, limiting the representation capacity and long-range dependencies that are crucial for deep neural networks [49].
-In 3D detection, long-range dependencies refer to connections between voxels within the same object and its immediate surroundings rather than between distant objects.
-Since LiDAR data is inherently sparse and occlusions are common, relying solely on local features makes it nearly impossible to detect objects like pedestrians with only a few sampled points.
-However, contextual cues, such as a crosswalk, can significantly aid detection, as even a few points above it may suggest the presence of a pedestrian.
-Therefore, long-range dependencies are essential for capturing contextual information.
-To address these limitations, the use of transformers has been proposed [46, 14, 47, 65, 43, 8, 48].
+Moreover, unlike the single-stride network design used in transformer-based detectors (e.g., [wang2023dsvt]), our SFMNet detector employs an encoder-based design.
+Voxel-based methods, which are the focus of this paper, transform irregular point clouds into regular voxel grids.
+This enables the use of convolutions, specifically sparse convolutions, that operate efficiently despite the presence of many empty voxels [choy20194d, sparseconv, second].
+While effective at capturing local structures, methods based on submanifold sparse convolution [graham2017submanifold] substantially limit the receptive field, thereby reducing the model’s ability to capture long-range dependencies, which are critical for detection [wang2018non].
+In 3D detection, long-range dependencies encompass relationships among voxels within an object as well as between the object and its immediate surroundings.
+Given the inherent sparsity and frequent occlusions in LiDAR data, relying only on local features makes detecting small objects, such as pedestrians, extremely difficult when only a few points are available.
+Contextual information, such as the presence of a crosswalk, can greatly aid detection, as even a few points above it may indicate a pedestrian’s presence, highlighting the importance of long-range dependencies.
+To address these limitations, the use of transformers has been proposed [chen2023focalformer3d, fan2022embracing, sun2022swformer, vaswani2017attention, wang2023dsvt, wang2023unitr, zhou2022centerformer].
 With self-attention (SA) as their core mechanism, transformers can capture global relationships, increasing the detector representation capacity.
 Despite its benefits, the quadratic complexity of SA is a significant bottleneck, particularly in autonomous vehicle scenarios where LiDAR scenes span hundreds of meters and contain hundreds of thousands of points.
-On the other hand, convolutional neural networks have recently made a comeback, primarily in 2D, through the use of large kernels [12, 11, 25].
-Large-kernel convolution-based methods improve representation capacity by increasing the receptive field and offering greater robustness compared to transformers [4].
-However, this approach is often impractical for 3D scenes, as the number of parameters grows cubically, necessitating techniques like sophisticated kernel weight partitioning with shared weights [6], data-driven kernel generation [27], or weight pruning [16].
+We propose a ’compromise’ between the two approaches by using large kernels in convolutional neural networks.
+This strategy has demonstrated efficiency, primarily in 2D [ding2022scaling, ding2024unireplknet, liu2022more].
+Large-kernel methods improve representation capacity by increasing the receptive field and offering greater robustness compared to transformers [chen2024revealing].
+However, this approach is often impractical for 3D scenes, as the number of parameters grows cubically, necessitating techniques like sophisticated kernel weight partitioning with shared weights [chen2023largekernel3d], data-driven kernel generation [lu2023link], or weight pruning [feng2024lsk3dnet].
 Although these methods reduce parameter bottlenecks and enlarge the effective receptive field, they offer minimal accuracy improvements and are less efficient than traditional sparse convolutions.
@@
 The core idea is to efficiently extract local contexts using sparse convolutions in a hierarchical structure, enabling the design to capture global relationships.
-Limiting attention to local context windows has been shown to reduce the complexity bottleneck [26].
-However, due to the sparse nature of LiDAR points, the number of occupied voxels in each local window can vary significantly, making efficient local-based attention non-trivial [47].
+Limiting attention to local context windows has been shown to reduce the complexity bottleneck [liu2021swin].
+However, due to the sparse nature of LiDAR points, the number of occupied voxels in each local window can vary significantly, making efficient local-based attention non-trivial [wang2023dsvt].
 To efficiently extract context windows, we design our SFM module with stacked submanifold sparse convolutions, using small kernels and varying dilation rates.
 This design enables hierarchical context aggregation, progressively expanding the receptive field at each level.
-Sparse convolutions capture local structures efficiently, while the hierarchical design and increased dilation rates allow for learning sparse long-range dependencies.
+Sparse convolutions capture local structures efficiently, while the hierarchical design and increased dilations allow for learning sparse long-range dependencies.
 We validate the effectiveness of our detector on three well-known autonomous driving datasets.
-Our results demonstrate that the network advances the state of the art in long-range detection (Argoverse2 [50]), for which it is designed.
-On mid- to short-range datasets, namely Waymo Open [41] and nuScenes [2], it performs comparably or slightly better than previous state-of-the-art detectors.
+Our results demonstrate that the network advances the state-of-the-art in long-range detection (Argoverse2 [wilson2023argoverse]), for which it is designed.
+On mid- to short-range datasets, Waymo Open [waymo] and nuScenes [nuscenes], it performs comparably or slightly better than previous state-of-the-art detectors.
 Thus, the main contributions are twofold:
-  • We propose a novel sparse focal modulation (SFM) module.
-It enables long-term dependencies, similar to transformers, while maintaining efficiency like sparse convolutions.
-This is achieved by hierarchically aggregating contextual information from sparse inputs.
-The SFM module outperforms traditional attention-based methods by reducing computational overhead and efficiently processing sparse voxels of varying densities within context windows.
+  • We propose a novel sparse focal modulation (SFM) module that enables long-range dependencies like transformers while maintaining the efficiency of sparse convolutions.
+It hierarchically aggregates contextual information from sparse inputs, outperforming traditional attention methods by reducing computational overhead and efficiently handling sparse voxels of varying densities.
   • We introduce SFMNet, a 3D object detector built on top of the SFM module, which effectively handles large point cloud scenes.
@@
 3D detection on point clouds.
-3D detection methods can be broadly categorized into point-based methods and voxel-based methods.
-Point-based methods use PointNet [32, 33]-like architectures to extract geometric features directly from raw point clouds [32, 33, 34, 35, 36, 38, 39, 54].
+3D detection methods are broadly categorized into point-based methods and voxel-based methods.
+Point-based methods use PointNet [qi2017pointnet, qi2017pointnet++]-like architectures to extract geometric features directly from raw point clouds [qi2017pointnet, qi2017pointnet++, qi2018frustum, pointrcnn, pvrcnn, pvrcnn++, shi2020point, yang20203dssd, wang2023sat].
 However, in large-scale point cloud scenes typical of outdoor LiDAR datasets in autonomous driving, these methods suffer from lower accuracy and high processing times due to point sampling and neighbor searching.
-In contrast, voxel-based methods convert irregular point clouds into regular voxel grids and apply sparse convolutions on non-empty voxels [63, 5, 10, 23, 37, 52, 55, 56, 61, 62, 7, 59, 58, 40].
+In contrast, voxel-based methods convert irregular point clouds into regular voxel grids and apply sparse convolutions on non-empty voxels [zhou2018voxelnet, chen2019fast, voxelrcnn, pointpillars, parta2, second, ye2020hvnet, yin2021center, zheng2021cia, zheng2021se, chen2023voxelnext, zhang2024hednet, zhang2024safdnet, shrout2023gravos, liu2024sparsedet].
 While efficient for processing large-scale data, these methods rely on submanifold sparse convolution, which might limit network representation capacity by prioritizing efficiency over capturing long-range dependencies.
@@
 Recent 3D detection methods capture long-range dependencies using large-kernel sparse convolutions or transformers.
-Large-kernel convolution approaches expand the receptive field by increasing kernel sizes, a technique that has shown accuracy gains in 2D detection [12, 11, 25].
+Large-kernel convolution approaches expand the receptive field by increasing kernel sizes, a technique that has shown accuracy gains in 2D detection [ding2024unireplknet, ding2022scaling, liu2022more].
 However, directly applying large kernels to 3D scenes incurs a cubically increasing computational overhead.
-To address this, LargeKernel3D [6] introduces a spatial partitioning convolution with shared weights among neighboring voxels.
-LinK [27] further optimizes efficiency by generating sparse kernels dynamically, assigning weights only to non-empty voxels rather than maintaining a dense kernel matrix.
-Similarly, LSK3DNet [16] reduces computation by pruning redundant weights across spatial and channel dimensions.
+To address this, LargeKernel3D [chen2023largekernel3d] introduces a spatial partitioning convolution with shared weights among neighboring voxels.
+LinK [lu2023link] further optimizes efficiency by generating sparse kernels dynamically, assigning weights only to non-empty voxels rather than maintaining a dense kernel matrix.
+Similarly, LSK3DNet [feng2024lsk3dnet] reduces computation by pruning redundant weights across spatial and channel dimensions.
 While these methods alleviate some computational demands, they still face performance challenges in large-scale 3D data.
 Transformer-based detectors, by contrast, use self-attention to capture long-range dependencies, enabling global token interactions.
-However, applying attention in high-resolution 3D scenes with many tokens is computationally prohibitive [65, 8, 48].
-To mitigate this, VoTr [30] implements local and dilated attention, focusing on sparse voxel interactions.
-SWFormer [43] utilizes a window-based transformer with a bucketing scheme to manage windows with varying numbers of non-empty voxels.
-SST [14] refines token processing by grouping tokens regionally, applying attention within each group, and shifting regions to capture structured interactions.
-DSVT [47] sequentially applies windowed self-attention along the X and Y axes to capture dependencies with reduced complexity.
-Although they improve long-range relationships, they still face quadratic scaling issues with context window tokens due to self-attention.
+However, applying attention in high-resolution 3D scenes with many tokens is computationally prohibitive [zhou2022centerformer, chen2023focalformer3d, wang2023unitr].
+To mitigate this, VoTr [mao2021voxel] implements local and dilated attention, focusing on sparse voxel interactions.
+SWFormer [sun2022swformer] utilizes a window-based transformer with a bucketing scheme to manage windows of varying non-empty voxels.
+SST [fan2022embracing] refines token processing by grouping tokens regionally, applying attention within each group, and shifting regions to capture structured interactions.
+DSVT [wang2023dsvt] sequentially applies windowed self-attention along the X and Y axes to capture dependencies with reduced complexity.
+Although they improve long-range relationships, they still face quadratic scaling issues with context-window tokens due to self-attention.
 We propose a network that combines the strengths of sparse convolutions and transformers.
@@
 They commonly leverage sparse convolutions for efficient learning in sparse environments.
-There are two main types of sparse convolutions: regular sparse convolution [18] and submanifold sparse convolution [19].
+There are two main types of sparse convolutions: regular [graham2015sparse] and submanifold [graham2017submanifold].
 Regular sparse convolutions expand features into (not necessarily occupied) neighboring voxels, enabling information exchange between spatially disconnected regions, but increasing feature map density.
-Due to their expansive nature, they are typically used for downsampling and upsampling.
-Submanifold sparse convolutions, on the other hand, operate strictly on occupied voxel features, ensuring that the output retains the same sparsity as the input, thereby facilitating efficient learning of local structures.
-However, due to their limited kernel size, they lack long-range dependencies, which are useful for object detection [49].
+Their expansive nature makes them suitable for downsampling and upsampling.
+Submanifold sparse convolutions operate only on occupied voxels, ensuring that the output retains the same sparsity as the input, thereby facilitating efficient learning of local structures.
+However, due to their limited kernel size, they lack long-range dependencies, which are useful for object detection [wang2018non].
 We aim to design an efficient detector built from sparse convolutions capable of capturing both short- and long-range dependencies.
-A straightforward approach to modeling global relationships would be to use transformers, which have shown great success in the 2D domain.
-However, applying standard transformers to sparse point clouds is challenging: (1) With numerous occupied voxels in each scene, standard self-attention suffers from massive query-key and query-value interactions, leading to a high computational overhead or suboptimal accuracy due to subsampling [31, 60].
-(2) Due to the sparse nature of LiDAR point clouds, the number of occupied voxels in each local region can vary significantly, leading to inefficient batch processing [14, 47].
+While transformers are effective for modeling global relationships in 2D, applying them to sparse point clouds is challenging: (1) With numerous occupied voxels in each scene, standard self-attention suffers from massive query-key and query-value interactions, leading to a high computational overhead or suboptimal accuracy due to subsampling [pan20213d, zhao2021point].
+(2) The varying voxel density across regions hinders efficient batching [fan2022embracing, wang2023dsvt].
 To address both the computational overhead and the varying number of occupied voxels within a context window, we propose replacing self-attention with Sparse Focal Modulation (SFM) as the token mixer in the transformer.
@@
 Additionally, sparse convolutions were designed to efficiently handle varying numbers of occupied voxels within each local region and can facilitate effective context extraction.
-The context extraction and interaction phases of our proposed SFM module are illustrated in Fig.
-2, in comparison to naive local-based attention.
-Local-based attention involves a context extraction phase for each query token, followed by query-key and query-value interactions.
-The context extraction phase requires identifying all non-empty voxels within the context window, sampling or padding them to a fixed maximum number of neighbors, and concatenating their features for batch processing.
-In high-resolution scenes with hundreds of thousands of voxels, this process quickly becomes a bottleneck.
-Moreover, the query-key-value interactions result in quadratic complexity relative to the number of tokens within the context window.
-In contrast, SFM modulates a query’s interactions by extracting multiple levels of context using sparse convolutions, focusing on short-range structures at lower levels and long-range dependencies at higher levels.
-The sparse convolutions are designed to handle varying numbers of non-empty voxels efficiently within different context windows, enabling effective batch processing.
-By using (sparse) convolutions with different receptive fields for context extraction, the interaction phase is reduced to only the number of extracted context levels.
 Our SFM module is described in Section 3.1.
+Fig.
+2 illustrates its context extraction and interaction phases compared to naive local attention.
+Local attention involves extracting a context for each query token, then performing query-key and query-value interactions.
+This requires identifying all non-empty voxels in the context window, sampling or padding them to a fixed maximum, and concatenating features for batch processing.
+In high-resolution scenes with hundreds of thousands of voxels, this quickly becomes a bottleneck, and query-key-value interactions incur quadratic complexity with the number of tokens per window.
+In contrast, SFM modulates query interactions by extracting multiple context levels via sparse convolutions, capturing short-range structures at lower levels and long-range dependencies at higher levels.
+These sparse convolutions efficiently handle varying numbers of non-empty voxels within context windows, enabling effective batch processing.
+By using convolutions with different receptive fields for context extraction, the interaction phase is reduced to only the number of context levels.
 The effectiveness of our proposed SFM is demonstrated through a novel sparse detector, called SFMNet, introduced in Section 3.2.
@@
 Focal Modulation (FM) is an alternative token mixer to self-attention, designed to limit token interactions.
-It has been applied in 2D tasks, such as classification, segmentation, and detection [53, 28].
+It has been applied in 2D tasks like classification, segmentation, and detection [yang2022focal, ma2024efficient].
 Extending FM to 3D faces two primary challenges.
-First, 3D representations are inherently sparse.
-The voxel count, including both empty and non-empty voxels, scales cubically.
-For example, in Argoverse2 [50], a typical scene covers an area of $400\times 400\times 8$ meters, resulting in approximately $640$ million voxels at a resolution of $0.1\times 0.1\times 0.2$ meters.
+First, 3D representations are inherently sparse, with voxel count scaling cubically.
+For instance, a typical Argoverse2 [wilson2023argoverse] scene spans $400\times 400\times 8$ meters, yielding approximately $640$ million voxels at $0.1\times 0.1\times 0.2$ m resolution.
 With only about 100,000 sampled points, this results in a high percentage of empty voxels.
-The second challenge is that 2D FM typically relies on large kernels to capture global context by using a stack of depth-wise convolutional layers with progressively larger kernels to capture visual contexts at various scales.
-However, in 3D object detection, sparse depth-wise convolutions have shown no advantage over spatial convolutions and are memory-intensive [6].
+Second, 2D FM typically captures global context using stacked depth-wise convolutions with progressively larger kernels to model multi-scale visual features.
+However, in 3D detection, sparse depth-wise convolutions offer no benefit over spatial ones and are memory-intensive [chen2023largekernel3d].
 Furthermore, applying large kernels in 3D scenes leads to a cubic increase in parameter count, substantially raising computational costs.
-We propose a new Sparse Focal Modulation (SFM) module that addresses these challenges.
-It captures both short- and long-range contexts while adaptively exchanging information with each token query.
+We propose a new Sparse Focal Modulation (SFM) module that addresses these challenges and captures both short- and long-range contexts while adaptively exchanging information with each token query.
 Unlike the conventional approach of using large kernels to expand the context window densely, we expand the context window sparsely with small kernels and increased dilation rates.
 We reason that queries related to small objects often rely on local structures and benefit from compact, focused context windows.
-In contrast, queries for large objects (e.g., partially occluded vehicles or roads) require broader context windows, but not necessarily dense windows, given the inherent sparsity of LiDAR point clouds.
+In contrast, queries for large objects (e.g., partially occluded vehicles or roads) require broader but not necessarily dense context windows, given the sparsity of LiDAR point clouds.
 Hence, our context aggregation is sparse.
+[…]
+Figure 3: Sparse Focal Modulation (SFM).
+SFM leverages sparse convolutions to aggregate multi-level contexts and modulate them with query features, efficiently capturing both local and global dependencies.
 Fig.
 3 illustrates our SFM module, which modulates each non-empty voxel feature ($x$) with surrounding information to produce a spatially aware feature ($z$).
-Its core component, the Context Aggregation (CA) block, captures relevant context for each query token by: 1) efficiently extracting contexts hierarchically, from local to global regions (e.g., red, green, and blue regions in Fig.
+Its core component, the Context Aggregation (CA) block, captures context for each query token by: 1) efficiently extracting contexts hierarchically, from local to global regions (e.g., red, green, and blue regions in Fig.
 2); and 2) adaptively aggregating context features at each level using a gating mechanism to provide meaningful context for each query.
-The resulting aggregated context of the CA block, ctx, is modulated with the corresponding query $q$ to generate the feature $z$.
-[…]
-Figure 3: Sparse Focal Modulation (SFM).
-Given non-empty voxel features $x$, the query $q$ and initial focal features $f^{0}$ are generated through linear layers.
-$f^{0}$ is fed into the Context Aggregation module, which aggregates multi-level contexts by stacking submanifold sparse convolutions (SubMConv) with small kernel sizes ($k$) and varying dilation rates ($d$), each followed by a GeLU activation.
-Context aggregation is achieved through a gated mechanism that adaptively combines features from each level ($f^{\ell}$) based on their relevance.
-The final context (ctx) is mapped to the query space using a linear layer $h(\cdot)$ and then modulated with the query $q$ to produce spatially-aware features $z$.
-This design captures both local and global dependencies while maintaining computational efficiency.
+The aggregated context of the CA block, ctx, is modulated with the corresponding query $q$ to generate the feature $z$.
 Submanifold sparse convolutions efficiently handle voxel sparsity, as they operate strictly on non-empty voxels.
@@
 $f^{\ell}=\text{GeLU}(\text{SubMConv}(f^{\ell-1};k^{\ell},d^{\ell})),$ (1)
-where $\text{SubMConv}(\cdot;k^{\ell},d^{\ell})$ denotes a submanifold sparse convolution with kernel size $k^{\ell}$ and dilation rate $d^{\ell}$, and GeLU is an activation function [21].
+where $\text{SubMConv}(\cdot;k^{\ell},d^{\ell})$ denotes a submanifold sparse convolution with kernel size $k^{\ell}$ and dilation rate $d^{\ell}$, and GeLU is an activation function [hendrycks2016gaussian].
 The hierarchical convolutions effectively capture multiple levels of focal abstraction, with the effective receptive field at level $\ell$ given by:
 $r^{\ell}=1+\sum_{i=1}^{\ell}(k^{i}-1)\cdot d^{i}.$ (2)
-As for the second key principle, when aggregating context features, we note that not all queries benefit equally from short-range or long-range contexts.
+When aggregating context features at each level, we note that not all queries benefit equally from short-range or long-range contexts.
 Specifically, queries for small objects are more likely to gain from the short-range context available at lower focal levels, while larger objects or background regions tend to benefit more from the long-range context captured at higher focal levels.
@@
 where $g^{\ell}\in\mathbb{R}^{N\times 1}$ represents the gate weight for focal level $\ell$, and $h(\cdot)$ is a linear layer that projects the output context into the query space.
-Finally, we modulate the focal contexts to the queries through element-wise multiplication:
+Finally, we modulate the focal contexts (ctx) to the queries ($q$) through element-wise multiplication, which enables each query to learn its relevant context:
 $z=q\cdot\text{ctx}\in\mathbb{R}^{N\times C}.$ (4)
-This enables each query to learn its relevant context.
-In summary, our module addresses the aforementioned challenges.
-By stacking sparse convolutions with small kernels and progressively increasing dilation, we effectively aggregate multi-level contexts.
-Efficiently handling voxel sparsity while modulating a query with its aggregated context maintains token interactions at linear complexity.
-Method mAPH Vehicle Pedestrian Cyclist
-Baseline 66.9 64.4 67.0 69.1
-SFMNet 69.5 68.1 69.2 71.1
-Table 1: The utility of our SFM module. Incorporating a single SFM module into each backbone (3D and 2D) within CenterPoint improves L2 mAPH by $2.6\%$ on a $20\%$ Waymo Open subset.
-To demonstrate the utility of our proposed module, we incorporated a single SFM module into each backbone (3D and 2D) within the CenterPoint [56] architecture, without otherwise altering the structure.
-As shown in Table 1, this addition yielded a $2.6\%$ improvement in the Level-2 mAPH on a subset of the Waymo Open dataset [41].
-Next, we will introduce our architecture that leverages our module.
+In summary, to address voxel sparsity in 3D space, we propose Sparse Focal Modulation (SFM), which expands the context window efficiently using submanifold sparse convolutions with small kernels and increasing dilation.
+To handle varying contextual needs, the Context Aggregation (CA) block employs a gating mechanism to adaptively aggregate multi-scale features, allowing each query to focus on the most relevant context for accurate 3D detection.
+To demonstrate its utility, we incorporated a single SFM module into each backbone (3D and 2D) within the CenterPoint [yin2021center] architecture, without otherwise altering the structure.
+This has already yielded a $2.6\%$ improvement in Level-2 mAPH on a $20\%$ subset of the Waymo Open Dataset [waymo].
 3.2 SFMNet
-Building on our sparse focal modulation (SFM) module, we present SFMNet, a simple and efficient sparse detector for 3D object detection.
+Building on our sparse focal modulation (SFM) module, we present SFMNet, a simple and efficient sparse 3D object detector.
 It enables learning long-range dependencies that convolution-based backbones lack, while maintaining sparse convolution efficiency.
 Our architecture is built on two core principles.
-First, increasing the effective receptive field (ERF) in mid- and high-level layers enhances the network’s capacity to capture broader spatial contexts.
-This is achieved by incorporating our proposed SFM module.
-Second, once a reasonable receptive field is achieved, scaling a model with small kernels becomes more efficient due to better parameter utilization.
-This aligns with [12]’s observation that beyond a certain point, adding large kernels is less efficient than adding small kernels.
-In our setup, SFM blocks can be seen as analogous to large kernels, while SRB blocks correspond to small kernels.
-Therefore, we include more SRB blocks than SFM blocks (see Section 4.2).
+First, increasing the effective receptive field (ERF) in mid- and high-level layers enhances the network’s ability to capture broader spatial context—this is primarily achieved using a single SFM block.
+Second, once a sufficient receptive field is established, adding depth with SRB blocks becomes more efficient, as they act like small kernels and offer better parameter utilization.
+In practice, we find that one SFM block combined with one SRB block is sufficient to meet our design goals.
+When further deepening the network, we prioritize adding more SRB blocks rather than additional SFM blocks (see Section 4.2).
+This design choice aligns with the observation from [ding2024unireplknet] that, after the receptive field is adequately covered, small kernels become more efficient than large ones—an analogy that maps well to our use of SFM (large kernel equivalent) and SRB (small kernel equivalent) blocks.
 […]
@@
 4(a) illustrates the SFMNet architecture.
-The input point cloud is voxelized, and features are extracted using a dynamic voxel feature extraction (VFE) method [64].
+The input point cloud is voxelized, and features are extracted using a dynamic voxel feature extraction (VFE) method [zhou2020end].
 These features are then processed by our 3D and 2D sparse backbones.
@@
 The higher the stage, the more SRB blocks per SFM block should be added to scale the model.
-The SFM block utilizes our SFM module as the token mixer within a standard MetaFormer-like design [57], as shown in Fig.
-4(c).
+The SFM block employs our SFM module as the token mixer within a standard MetaFormer-like design [yu2022metaformer], as illustrated in Fig.
+4(c), where non-empty voxels serve as the tokens.
 Formally, given an input feature map $x\in\mathbb{R}^{N\times C}$ and its corresponding focal-modulated feature $z\in\mathbb{R}^{N\times C}$ generated by the SFM module, as described in Section 3.1, the SFM block is formulated as:
-$y^{\prime}$ $=\text{LN}(z)+x$ (5)
-$y$ $=\text{LN}(\text{MLP}(y^{\prime}))+y^{\prime},$ (6)
-where LN denotes LayerNorm, and the multilayer perceptron (MLP) contains a single hidden layer.
-The 3D sparse features are then compressed into 2D sparse bird’s-eye view (BEV) features, following the approach in [58], and subsequently fed into a 2D backbone.
-Similar to the 3D backbone, the 2D backbone also consists of an SFM block followed by SRBs but employs 2D sparse convolutions instead of 3D.
+$y^{\prime}=\text{LN}(z)+x,\quad y=\text{LN}(\text{MLP}(y^{\prime}))+y^{\prime},$ (5)
+where LN denotes LayerNorm, the multilayer perceptron (MLP) has a single hidden layer, and $y$ is the refined feature from focal sparse modulation.
+The 3D sparse features are then compressed into 2D bird’s-eye view (BEV) features, as in [zhang2024safdnet], and passed to a 2D backbone.
+Like the 3D backbone, the 2D backbone includes an SFM block followed by SRBs, but uses 2D sparse convolutions.
 4 Experiments
 Datasets and metrics.
-To validate our approach’s effectiveness, we conducted experiments on three popular datasets: Argoverse2 [50], Waymo Open [41], and nuScenes [2].
-The detection range of these datasets varies, with long-range detection in Argoverse2 (up to ±$\pm 200$ meters), mid-range in Waymo Open (up to ±$\pm 75$ meters), and short-range in nuScenes (up to ±$\pm 54$ meters).
+To validate our approach’s effectiveness, we conducted experiments on three popular datasets: Argoverse2 [wilson2023argoverse], Waymo Open [waymo], and nuScenes [nuscenes].
+The detection range of these datasets varies, with long-range detection in Argoverse2 (up to $\pm 200$ meters), mid-range in Waymo Open (up to $\pm 75$ meters), and short-range in nuScenes (up to $\pm 54$ meters).
 We use the standard evaluation metrics for each dataset: mAP for Argoverse2, mAP and mAP weighted by heading accuracy (mAPH) for Waymo Open, and mAP and nuScenes detection score (NDS) for nuScenes.
@@
 W-Ride
-CenterPoint [56] 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
-VoxelNeXt [7] 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
-HEDNet [59] 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
-SAFDNet [58] 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
-SFMNet (Ours) 40.4 79.0 49.5 72.3 49.3 46.1 65.1 73.8 55.1 53.2 62.8 54.2 30.2 44.8 26.7 57.9 31.7 20.9 8.3 33.4 44.4 21.6 0.0 22.1 3.1 33.1 12.1
-Table 2: Results on the Argoverse2 validation set (long-range, up to ±$\pm 200$ meters). The long-range detection dataset highlights our detector’s ability to capture long-range dependencies.
+CenterPoint 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
+VoxelNeXt 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
+HEDNet 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
+SAFDNet 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
+SFMNet 40.4 79.0 49.5 72.3 49.3 46.1 65.1 73.8 55.1 53.2 62.8 54.2 30.2 44.8 26.7 57.9 31.7 20.9 8.3 33.4 44.4 21.6 0.0 22.1 3.1 33.1 12.1
+Table 1: Results on the Argoverse2 validation set (long-range, up to $\pm 200$ meters). The long-range detection dataset highlights our detector’s ability to capture long-range dependencies.
 Design Method mAP/mAPH Vehicle AP/APH Pedestrian AP/APH Cyclist AP/APH
 L1 L2 L1 L2 L1 L2 L1 L2
-Transformers SST [14] [CVPR’22] 74.5/71.0 67.8/64.6 74.2/73.8 65.5/65.1 78.7/69.6 70.0/61.7 70.7/69.6 68.0/66.9
-TransFusion-L [1] [CVPR’22] -/- -/64.9 -/- -/65.1 -/- -/63.7 -/- -/65.9
-SWFormer [43] [ECCV’22] -/- -/- 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 -/- -/-
-CenterFormer [65] [ECCV’22] 75.6/73.2 71.4/69.1 75.0/74.4 69.9/69.4 78.0/72.4 73.1/67.7 73.8/72.7 71.3/70.2
-DSVT-Voxel [47] [CVPR’23] 80.3/78.2 74.0/72.1 79.7/79.3 71.4/71.0 83.7/78.9 76.1/71.5 77.5/76.5 74.6/73.7
-PTv3 [51] [CVPR’24] -/- 73.0/70.5 -/- 71.2/70.8 -/- 76.3/70.4 -/- 71.5/70.4
-Large kernels LargeKernel3D [6] [CVPR’23] -/- -/- 78.1/77.6 69.8/69.4 -/- -/- -/- -/-
-LinK [27] [CVPR’23] (6 epochs) -/- -/64.6 -/- -/65.0 -/- -/60.4 -/- -/68.4
-Small kernels SECOND [52] [Sensors’18] 67.2/63.1 61.0/57.2 72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0
-PointPillars [23] [CVPR’19] 69.0/63.5 62.8/57.8 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9
-Part-A2 [37] [TPAMI’20] 73.6/70.3 66.9/63.8 77.1/76.5 68.5/68.0 75.2/66.9 66.2/58.6 68.6/67.4 66.1/64.9
-PV-RCNN [36] [CVPR’20] 76.2/73.6 69.6/67.2 78.0/77.5 69.4/69.0 79.2/73.0 70.4/64.7 71.5/70.3 69.0/67.8
-CenterPoint [56] [CVPR’21] 75.9/73.5 69.8/67.6 76.6/76.0 68.9/68.4 79.0/73.4 71.0/65.8 72.1/71.0 69.5/68.5
-AFDetV2 [22] [AAAI’22] 77.2/74.8 71.0/68.8 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1
-FSD [15] [NeurIPS’22] 79.6/77.4 72.9/70.8 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3
-PV-RCNN++ [38] [IJCV’23] 78.1/75.9 71.7/69.5 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2
-VoxelNeXt [7] [CVPR’23] 78.6/76.3 72.2/70.1 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2
-HEDNet [59] [NeurIPS’23] 81.4/79.4 75.3/73.4 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9
-SAFDNet [58] [CVPR’24] 81.8/79.9 75.7/73.9 80.6/80.1 72.7/72.3 84.7/80.4 77.3/73.1 80.0/79.0 77.2.76.2
+Transformers SST [fan2022embracing] 74.5/71.0 67.8/64.6 74.2/73.8 65.5/65.1 78.7/69.6 70.0/61.7 70.7/69.6 68.0/66.9
+TransFusion-L [bai2022transfusion] -/- -/64.9 -/- -/65.1 -/- -/63.7 -/- -/65.9
+SWFormer [sun2022swformer] -/- -/- 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 -/- -/-
+CenterFormer [zhou2022centerformer] 75.6/73.2 71.4/69.1 75.0/74.4 69.9/69.4 78.0/72.4 73.1/67.7 73.8/72.7 71.3/70.2
+DSVT-Voxel [wang2023dsvt] 80.3/78.2 74.0/72.1 79.7/79.3 71.4/71.0 83.7/78.9 76.1/71.5 77.5/76.5 74.6/73.7
+PTv3 [wu2024point] -/- 73.0/70.5 -/- 71.2/70.8 -/- 76.3/70.4 -/- 71.5/70.4
+Large kernels LargeKernel3D [chen2023largekernel3d] -/- -/- 78.1/77.6 69.8/69.4 -/- -/- -/- -/-
+LinK [lu2023link] (6 epochs) -/- -/64.6 -/- -/65.0 -/- -/60.4 -/- -/68.4
+Small kernels SECOND [second] 67.2/63.1 61.0/57.2 72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0
+PointPillars [pointpillars] 69.0/63.5 62.8/57.8 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9
+Part-A2 [parta2] 73.6/70.3 66.9/63.8 77.1/76.5 68.5/68.0 75.2/66.9 66.2/58.6 68.6/67.4 66.1/64.9
+PV-RCNN [pvrcnn] 76.2/73.6 69.6/67.2 78.0/77.5 69.4/69.0 79.2/73.0 70.4/64.7 71.5/70.3 69.0/67.8
+CenterPoint [yin2021center] 75.9/73.5 69.8/67.6 76.6/76.0 68.9/68.4 79.0/73.4 71.0/65.8 72.1/71.0 69.5/68.5
+AFDetV2 [hu2022afdetv2] 77.2/74.8 71.0/68.8 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1
+FSD [fan2022fully] 79.6/77.4 72.9/70.8 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3
+PV-RCNN++ [pvrcnn++] 78.1/75.9 71.7/69.5 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2
+VoxelNeXt [chen2023voxelnext] 78.6/76.3 72.2/70.1 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2
+HEDNet [zhang2024hednet] 81.4/79.4 75.3/73.4 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9
+SAFDNet [zhang2024safdnet] 81.8/79.9 75.7/73.9 80.6/80.1 72.7/72.3 84.7/80.4 77.3/73.1 80.0/79.0 77.2.76.2
 SFMNet (ours) 81.8/79.8 75.7/73.8 80.5/80.0 72.6/72.2 85.0/80.6 77.4/73.2 79.9/78.9 77.0/76.1
-Table 3: Results on the Waymo Open validation set (mid-range, up to ±$\pm 75$ meters).
-Our SFMNet outperforms transformer- and large-kernel-based detectors, while achieving results comparable to the state-of-the-art small-kernel detector [58].
+Table 2: Results on the Waymo Open validation set (mid-range, up to $\pm 75$ meters).
+Our SFMNet outperforms transformer- and large-kernel-based detectors, while achieving results comparable to the state-of-the-art small-kernel detector [zhang2024safdnet].
+Design Method NDS mAP Car Truck Bus Trailer C.V. Pedestrian Motor Bike Cone Barrier
+Transformers TransFusion-L [bai2022transfusion] 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
+DSVT-Voxel [wang2023dsvt] 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.8 70.9
+Large kernels LargeKernel3D [chen2023largekernel3d] 69.1 63.9 85.1 60.1 72.6 41.4 24.3 85.6 70.8 59.2 72.3 67.7
+LinK [lu2023link] 69.5 63.6 - - - - - - - - - -
+Small kernels CenterPoint [yin2021center] 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
+VoxelNeXt [chen2023voxelnext] 66.7 60.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
+PillarNeXt-L [li2023pillarnext] 68.4 62.2 85.0 57.4 67.6 35.6 20.6 86.8 68.6 53.1 77.3 69.7
+HEDNet [zhang2024hednet] 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
+SAFDNet [zhang2024safdnet] 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
+SFMNet (ours) 71.3 66.9 87.6 63.2 76.2 48.9 26.0 88.4 74.7 56.9 78.5 69.1
+Table 3: Results on the nuScenes validation set (short-range, up to $\pm 54$ meters).
+Our SFMNet outperforms all transformer- and large-kernel-based detectors, surpassing the previous SoTA small-kernel detector on the mAP metric and ranking second on the NDS metric.
 Implementations details.
@@
 Ablation studies used 6 epochs on Argoverse2.
-Further implementation details are in the supplemental materials.
+Additional implementation details are provided in the supplemental.
 4.1 Results
@@
 We expect our model to be particularly beneficial for long-range scenes.
-Indeed, as shown in Table 2, which compares our method to state-of-the-art sparse convolution-based detectors with small kernels, our detector achieves a $0.7\%$ mAP improvement over the previous best method [58], demonstrating its ability to capture long-range dependencies.
+Indeed, as shown in Table 3, which compares our method to state-of-the-art sparse convolution-based detectors with small kernels, our detector achieves a $0.7\%$ mAP improvement over the previous best method [zhang2024safdnet], demonstrating its ability to capture long-range dependencies.
 For mid-range detection on Waymo Open, Table 3 compares our method to state-of-the-art methods on the validation set, where ’-’ indicates unavailable results.
-Our detector outperforms both transformer-based and large kernel-based methods, achieving gains of $1.7\%$ and $3.3\%$ mAPH on Level-2 difficulty compared to the transformer-based detectors DSVT [47] and PTv3 [51], respectively, while also achieving comparable results to the best small kernel-based detector [58].
+Our detector outperforms both transformer-based and large kernel-based methods, achieving gains of $1.7\%$ and $3.3\%$ mAPH on Level-2 difficulty compared to the transformer-based detectors DSVT [wang2023dsvt] and PTv3 [wu2024point], respectively, while also achieving comparable results to the best small kernel-based detector [zhang2024safdnet].
 For the short-range detection dataset nuScenes,Table 3 shows that our method achieves SoTA results on the mAP metric and ranks second on the NDS metric, outperforming both large kernel and transformer-based methods.
@@
 Figure 5: Effective receptive field (ERF) of SFMNet.
-(a) A typical input scene with height color-coded in shades of gray.
-(b,c) The ERF for each ground-truth object, color-coded from blue (outside ERF) to red (high attention regions).
-Red and orange rectangles highlight zoomed-in views of two regions.
+(a) A typical input scene with height is shown in grayscale.
+(b,c) The ERF for each ground-truth object is color-coded from blue (outside ERF) to red (high-attention regions).
+Red and orange rectangles highlight zoomed-in regions.
 In the right (orange) zoom-in, four stationary cars and a van, plus two moving cars and a van, are shown.
 Without the SFM module (b), only a local portion of each vehicle is within the ERF.
-With the SFM module (c), the entire object is captured, providing more context and spatial understanding for accurate 3D detection.
+With SFM (c), the entire object is captured, providing more context and spatial understanding for accurate 3D detection.
 Qualitative results.
@@
 For each ground-truth object, we randomly select one voxel to serve as the query and calculate the ERF associated with it.
-As intended, the dependency coverage is regional, resulting in large regional receptive fields.
-Without the SFM module (b), the detector’s receptive field is limited to local areas around each query, capturing only fragments of the objects.
-For instance, in the left zoom-in (red rectangle) in Fig.
+Indeed, the dependency coverage leads to large regional receptive fields.
+Without the SFM module (b), the receptive field is limited to local areas around each query, capturing only fragments of the objects.
+For instance, in the left zoom-in in Fig.
 5(b), there are four vehicles.
@@
 This is reflected in the four vehicles being highlighted in green, yellow, and red colors.
-Design Method NDS mAP Car Truck Bus Trailer C.V. Pedestrian Motor Bike Cone Barrier
-Transformers TransFusion-L [1] [CVPR’22] 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
-DSVT-Voxel [47] [CVPR’23] 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.8 70.9
-Large kernels LargeKernel3D [6] [CVPR’23] 69.1 63.9 85.1 60.1 72.6 41.4 24.3 85.6 70.8 59.2 72.3 67.7
-LinK [27] [CVPR’23] 69.5 63.6 - - - - - - - - - -
-Small kernels CenterPoint [56] [CVPR’21] 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
-VoxelNeXt [7] [CVPR’23] 66.7 60.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
-PillarNeXt-L [24] [CVPR’23] 68.4 62.2 85.0 57.4 67.6 35.6 20.6 86.8 68.6 53.1 77.3 69.7
-HEDNet [59] [NeurIPS’23] 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
-SAFDNet [58] [CVPR’24] 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
-SFMNet (ours) 71.3 66.9 87.6 63.2 76.2 48.9 26.0 88.4 74.7 56.9 78.5 69.1
-Table 4: Results on the nuScenes validation set (short-range, up to ±$\pm 54$ meters).
-Our SFMNet outperforms all transformer- and large-kernel-based detectors, surpassing the previous SoTA small-kernel detector on the mAP metric and ranking second on the NDS metric.
-4.2 Ablation study
-Table 5 compares our SFM module to traditional global and local attention, as well as set attention [47], on the Argoverse2 dataset over six epochs.
-For this evaluation, we replaced an attention block in the 3D and 2D backbones with either a single SFM module or an alternative attention mechanism, demonstrating that our SFM achieves higher accuracy and efficiency.
-While global attention attains a high ERF, it quickly reaches memory limits ($80$ GB) due to the large voxel count.
-In contrast, local and set attention operate within a limited context window, with set attention outperforming local attention.
-However, our SFM module surpasses both in accuracy (mAP) and efficiency (memory usage) despite using a smaller context window.
-Method Window size(m) Memory(GB) mAP
+Method Window(m) Memory(GB) mAP
 Global-attention 400 OOM -
 Local-attention 3.3 44.06 34.8
-Set-attention (DSVT [47]) 3.6 44.21 35.5
+Set-attention ([wang2023dsvt]) 3.6 44.21 35.5
 SFM (ours) 1.9 32.76 36.8
 SFM (ours) 3.3 37.53 37.1
-Table 5: Comparison to attention mechanism.
-We replace our SFM module with various attention mechanisms and report mAP and per-GPU memory peak usage.
+Table 4: Comparison to attention.
 While global attention covers the entire ERF, it quickly exceeds memory limits.
-In contrast, Local-attention and Set-attention [47] restrict their focus to voxels within a local context window.
-Our SFM module outperforms both in accuracy and efficiency, even with a smaller context window.
-Table 6 examines the effect of varying the number of SFM blocks and SRBs.
-It shows that increasing the number of SRBs after the SFM block in Stage 4 (A) improves accuracy with a modest parameter increase.
-Similarly, adding more SFM block repetitions with SRBs (B) boosts accuracy but at a higher parameter cost.
-A balanced approach, with moderate repetitions and additional SRBs (C), achieves a good trade-off between accuracy and parameters.
-More ablation studies on SFM module placement and ERF size choices are provided in the supplemental materials.
+In contrast, Local-attention and Set-attention [wang2023dsvt] focus on voxels within a local context window.
+Our SFM module outperforms both in accuracy and efficiency, even with a smaller window.
 Stage 4 2D backbone #Param.(M) mAP
-#SRB Multiplier #SRB Multiplier
+#SRB Multi. #SRB Multi.
 A 2 1 4 1 3.84 36.7
@@
 C 4 2 4 2 6.16 38.9
-Table 6: Effect of SFM blocks and SRBs.
-Increasing SRBs after each SFM block in Stage 4 and the 2D backbone boosts accuracy.
-Adding more SFM block repetitions (Multiplier) further improves accuracy, though at the cost of more parameters.
-A balanced trade-off between accuracy and parameters is achieved in (C).
-Limitations. SFMNet relies on sparse convolutions with custom CUDA operations, making it less deployment-friendly than other deep learning frameworks (e.g., PyTorch).
+Table 5: Effect of SFM blocks & SRBs.
+Stacking SRBs after each SFM block improves accuracy (A).
+Increasing SFM repetitions boosts it further (B), but with more parameters.
+Configuration (C) offers a balanced trade-off.
+4.2 Ablation study
+Table 4 compares our SFM module to traditional global and local attention, as well as set attention [wang2023dsvt], on the Argoverse2 dataset over six epochs.
+For this evaluation, we replaced an attention block in the 3D and 2D backbones with either a single SFM module or an alternative attention mechanism, demonstrating that our SFM achieves higher accuracy and efficiency.
+While global attention attains a high ERF, it quickly reaches memory limits ($80$ GB) due to the large voxel count.
+In contrast, local and set attention operate within a limited context window, with set attention outperforming local attention.
+However, our SFM module surpasses both in accuracy (mAP) and efficiency (memory usage) despite using a smaller context window.
+Table 5 examines the effect of varying the number of SFM blocks and SRBs.
+Increasing the number of SRBs after the SFM block in Stage 4 (A) improves accuracy with a modest parameter increase.
+Similarly, adding more SFM block repetitions (Multipliers) with SRBs (B) boosts accuracy but at a higher parameter cost.
+A balanced approach, with moderate repetitions and additional SRBs (C), provides a good trade-off between accuracy and parameters.
+#levels dilation rates ERF(m) Memory(GB) mAP
+2 [1,3] 0.9 30.19 35.7
+3 [1,3,5] 1.9 32.76 36.8
+3 [1,5,9] 3.1 32.98 36.4
+4 [1,3,5,7] 3.3 37.53 37.1
+5 [1,3,5,7,9] 5.1 36.39 37.1
+Table 6: ERF effects.
+We evaluate different ERF sizes by increasing the number of context levels, each using a sparse convolution with kernel size $3$ and varying dilation.
+Four levels, yielding an ERF of $3.3$ meters, achieve the best accuracy.
+Table 6 investigates the impact of different context levels with varying ERF.
+Increasing levels improves accuracy but also raises memory usage.
+Beyond a point, more levels do not help, as the network tends to ignore overly long-range features.
+Relying only on high dilation rates without local contexts (third row) yields suboptimal performance, highlighting the need for short-range dependencies.
+Additional ablations on SFM placement are in the supplemental.
+Limitations.
+Our method is designed to capture long-range dependencies and enhance models that lack this capability.
+On datasets requiring long-range reasoning, it improves not only those models but also others with a similar objective.
+However, on short-range data, while it still benefits the former, its performance is on par with the latter.
 5 Conclusion
-This paper introduced a novel 3D Focal Modulation module, Sparse Focal Modulation (SFM), designed to effectively capture dependencies from short- to long-range.
+This paper introduced a novel 3D Focal Modulation module, SFM, designed to effectively capture dependencies from short- to long-range.
 It is based on a hierarchical submanifold sparse convolution design to efficiently handle data sparsity.
 It also presented SFMNet, a new sparse convolution-based 3D detector built on top of the SFM.
-It addresses the limitation of traditional (sparse) convolution detectors by significantly enlarging the effective receptive field.
-Experimental results demonstrate that SFMNet achieves state-of-the-art performance on long-range detection while delivering comparable or superior results on mid- to short-range datasets.
+It addresses the limitation of traditional sparse convolution detectors by significantly enlarging the effective receptive field.
+Experimental results demonstrate that SFMNet captures dependencies very well and achieves SoTA performance on multi-range datasets.
+[…]
+Supplementary Material
+6 Implementation details
+6.1 Network architecture
+For Argoverse2 and Waymo Open, we adopted the sparse version of CenterPoint [yin2021center] as the detection head, as used in [zhang2024safdnet].
+For nuScenes, we employed TransFusion [bai2022transfusion] as the detection head, in line with [zhang2024hednet, wang2023dsvt].
+Argoverse2.
+We begin by validating the effectiveness of our detector in the long-range detection of Argoverse2.
+The voxel sizes are set to $(0.1,0.1,0.2)$ meters.
+In stages 1 through 3, we employ $(0,1,1)$ SFM blocks, each followed by $(2,2,4)$ SRBs, respectively.
+In stage 4 and within the 2D backbone, we use $(4,2)$ SFM blocks, each followed by $(2,4)$ SRBs.
+For each SFM block, we set $L=4$ with kernel sizes of $(3,3,3,3)$ and dilation rates of $(1,3,5,7)$.
+Waymo Open.
+For the mid-range dataset, we set the voxel sizes to $(0.08,0.08,0.15)$ meters.
+In stages 1 through 3, we employ $(0,1,1)$ SFM blocks, each followed by $(2,2,4)$ SRBs, respectively.
+In stage 4 and within the 2D backbone, we use $(2,2)$ SFM blocks, each followed by $(6,6)$ SRBs.
+Since the Waymo scanning range is not as large as that of Argoverse2, we set $L=4$ with kernel sizes of $(3,5,3,5)$ and dilation rates of $(1,1,3,3)$ for each SFM block.
+This configuration results in an effective receptive field (ERF) of $2$ meters, compared to $3.3$ meters in Argoverse2.
+nuScenes.
+For the short-range detection dataset, we set the voxel sizes to $(0.075,0.075,0.2)$ meters and used the same configuration of SFM blocks and SRBs as in Waymo Open.
+Since nuScenes focuses on short-range detection (up to $\pm 54$ meters), it was shown in [zhang2024safdnet] that a sparse detection head offers no advantage over a dense one, as used in [zhang2024hednet].
+Consequently, we implemented the 2D backbone with dense convolutions instead of sparse ones.
+6.2 Training and inference schemes
+We implemented our approach using the OpenPCDet [openpcdet2020] framework for the Argoverse2 and Waymo Open datasets and the MMDetection3D [mmdet3d2020] framework for nuScenes, following the setup of state-of-the-art methods [wang2023dsvt, zhang2024hednet, zhang2024safdnet].
+Argoverse2.
+We follow the same training schemes as [yin2021center] to optimize the network using the Adam optimizer with a weight decay of $0.05$ and a one-cycle learning rate policy.
+The maximum learning rate is set to $5\times 10^{-3}$, and training is performed with a batch size of $32$ for $24$ epochs on $8$ NVIDIA L40 GPUs.
+Following [zhang2024safdnet], we apply ground-truth copy-paste data augmentation during training, disabling this augmentation in the final epoch as part of a fade strategy.
+For reporting frames per second (FPS) during inference, we use a single NVIDIA L40 GPU.
+Waymo Open.
+As in Argoverse, we follow the same training schemes as [yin2021center] to optimize the network using the Adam optimizer with a weight decay of $0.05$, a one-cycle learning rate policy, and a maximum learning rate of $5\times 10^{-3}$.
+Training is performed with a batch size of $32$ for $24$ epochs on $8$ NVIDIA A100 GPUs.
+Following [wang2023dsvt, zhang2024hednet, zhang2024safdnet], we apply ground-truth copy-paste data augmentation during training, disabling this augmentation in the final epoch as part of a fade strategy.
+During inference, and as described in [wang2023dsvt, zhang2024hednet, zhang2024safdnet], we use class-specific NMS with IoU thresholds of $0.7$, $0.6$, and $0.55$ for vehicles, pedestrians, and cyclists, respectively.
+nuScenes.
+We follow the training scheme adopted in [bai2022transfusion].
+The network is trained using the AdamW optimizer with a weight decay of $0.01$, a one-cycle learning rate policy, a maximum learning rate of $1\times 10^{-3}$, and a batch size of $32$ for $20$ epochs on $8$ NVIDIA A100 GPUs.
+Similarly to [bai2022transfusion, wang2023dsvt] we apply the fade strategy in the final $5$ epochs.
+Stage2 Stage3 Stage4 2D backbone mAP/mAPH
+L1 L2
+75.6/73.3 69.0/66.9
+76.4/74.2 70.1/67.9
+78.0/75.7 71.7/69.5
+77.5/75.2 71.2/69.0
+77.9/75.6 71.6/69.4
+Table 7: Comparison of SFM module placement.
+We integrate a single SFM module into various stages of the 3D backbone and the sole stage of the 2D backbone in our detector, evaluating on a $20\%$ subset of the Waymo Open dataset.
+The results indicate that focusing on stage $4$ of the 3D backbone and the 2D backbone is generally sufficient for optimal performance.
+7 Additional ablation studies
+To save computational resources, we conducted ablations on $20\%$ of the Waymo Open dataset [waymo], training for $12$ epochs on $8$ GPUs with a total batch size of $32$ and reporting results on the full validation set.
+We used a tiny SFMNet with a single SFM module in the 3D backbone and another in the 2D backbone, followed by $2$ and $4$ SRBs, respectively.
+To study SFM placement, we positioned the module at different backbone stages, and Table 7 shows that placing it in stage $4$ of the 3D backbone and in the 2D backbone is generally sufficient.